OpenStudy (anonymous):

Shelby currently has 43 Facebook friends. She is so popular that she adds 7 friends each week. a). Write an equation for the number of friends Shelby will have (f) after any number of week (w) b).Suppose 9 weeks have gone by. Use you equation to find the number of Facebook friends. c).Suppose Shelby has 155 friends. Set up and solve an equation to find the number of weeks that have gone by.

OpenStudy (anonymous):

Help please!

OpenStudy (anonymous):

a, f(w)=43+7w b, f(9)=43+7*9=106 c, 155=43+7x 7x=112 x=16

OpenStudy (anonymous):

Thanks :)